OpenStudy (anonymous):

Find the result when 30 is decreased by 85%.

OpenStudy (amistre64):

divide 30 up into 100 little pieces: 30/100 = 3/10 = .3 now how add 85 of those together; or simply mutliply it by 85 85 .3 ---- 25.5 30 - 25.5 = ?

OpenStudy (anonymous):

4.5?

OpenStudy (anonymous):

Change 85% to decimal form... .85 30x.85=25.5 then decrease the product from the last equation. 30-25.5= 4.5

OpenStudy (amistre64):

4.5 is good

OpenStudy (anonymous):

30*.85=25.5 30-25.5=4.5

OpenStudy (anonymous):

thanks